# Global dynamical structures from infinitesimal data

### v1

@@ Abstract
We provide a proof of Theorem 1 in Appendix I. The key implication of this result is that for sufficiently small $C_{G}$, the estimator $f$ will be away from 00 together with the distance $|y_{*}|$ to the attractor, in a quantifiable manner, allowing one to ascertain the proximity to it. We remark that the definition, and therefore the estimation and computation of $f$, requires exclusively linear operations on local data in the neighborhood of a point, and does not depend on the split of the coordinate system into $x$ and $y$, despite the fact that the formulation of our estimate does. By computing $f$ for all (or an appropriate subsample) of the trajectory data, and then filtering the $f$-sublevel sets of the full trajectory dataset (geometrically, a point cloud in $\mathbb{R}^{D}$), we can isolate the set of trajectory segments that lie on or quantifiably near any physically relevant attractor $\mathcal{A}$ in the global embedding (observation) space coordinates $\mathbb{R}^{D}$.
@@
We detail algorithms for computing $f(z)$ on measurement data in the Supplement. The computation is, at a high level, performed as follows. First we estimate a local Cartesian basis for $T_{z_{*}}\mathcal{M}\simeq T_{x_{*}}\mathcal{A}\oplus\mathbb{R}^{n-a}$ and compute the restriction of the local vector field to this subspace $v_{z_{*}}$. Local PCA, performed on the subsample of points lying within an open Euclidean ball of radius $r$ centered at $z_{*}$, provides a simple approach for computing a local basis of $T_{z_{*}}\mathcal{M}$ such that estimators for the tangent space and intrinsic dimension enjoy strong formal guarantees [41]. The local sample of the vector field $V_{z_{*}}:\mathbb{R}^{D}\rightarrow T\mathbb{R}^{D}$ can be approximated from finite differences. The local representation of the ($n$-dimensional) $v_{z_{*}}$ can then be found by orthogonal projection of the ($D$-dimensional) $V_{z_{*}}$ onto the vielbein basis. The Jacobian $H(z_{*})$ can then be estimated by centering the local vector field data and using linear estimation techniques; for experiments in the following sections, we use least squares regression with the standard MSE residual.
Figure 3: VERT distance estimator is sensitive to attractor hierarchies. A. Visualization of a trajectory $z_{i}(t)$ for a system with control parameters $\gamma_{\theta}=1$, $\gamma_{\rho}=2$, $\gamma_{\delta}=1$. From $\mathbf{H_{3}}$ (Table 1), we expect the trajectory to first reach the submanifold $\mathcal{C}$ (visualized in grey) before reaching the limit cycle $\mathcal{S}$. B. Representative plot of the distance estimator $f(z)$ (blue), and error coordinates for the intermediate attracting sets $\mathcal{P}$ ($e_{\delta}$, black dashed) and $\mathcal{C}$ ($e_{\rho}$, red dashed). The grey line indicates when the trajectory reaches $\mathcal{C}$, corresponding to a local minimum of $f(z)$. After the $\gamma_{\rho}$ control module has converged, the behavior of $f(z)$ is determined by the remaining $\gamma_{\delta}$ control module. The global minimum of $f(z)$ aligns with convergence to the attracting limit cycle $\mathcal{S}$ at the lowest level of the attractor hierarchy. C-D. We sweep combinations of the parameters $(\gamma_{\rho},\gamma_{\delta})$ to test sensitivity of $f(z)$ to intermediate attractors for systems without (C., $\gamma_{\theta}=0$) and with (D., $\gamma_{\theta}=1$) a limit cycle. Each column corresponds to an attracting submanifold hypothesis (c.f., Table 1). The MSE residual captures relative proximity of the identified ’template’ trajectory segments to the column’s submanifold, with lower values indicating closer proximity.
@@ Results
Detecting the presence of subprocesses evolving on distinct time scales is critical for model inference and interpretation of systems constructed by parallel composition of stable subsystems. We hypothesize that the ILG nature of the VERT pipeline affords sensitivity to both the global and intermediate attracting sets through a detectable change in the behavior of the distance estimator along a trajectory as it transitions between them.
We test this hypothesis using simulated trajectories from the A-HM (Eqs. 1 & 2) while systematically varying the control parameters $\gamma_{\rho}$ and $\gamma_{\delta}$, which determine the timescale of convergence for the two subsystems. We use VERT to predict the trajectory segments that lie on the globally invariant attracting set, and then estimate the proximity of these segments to each intermediate attracting submanifold outlined in Fig. 2 using the MSE projection residual, with lower values indicating closer alignment with the hypothesized submanifold. For the submanifolds indicated in the rightmost column, the first column indicates the $(\gamma_{\rho},\gamma_{\delta})$ parameter regime where the projection residual should be minimal to support the hypothesis. The results are summarized in Fig. 3 for A-HM equations with and without a globally attracting limit cycle ($\gamma_{\theta}=1$ and $\gamma_{\theta}=0$, respectively).
Evaluation of the attracting set predictions from VERT (Fig 3C-D) support the hypothesis that the behavior of the distance estimator along typical trajectories reflects the locally dominant (i.e., fastest) component of the vector field. Visualizing the evolution of the distance estimator over a representative trajectory provides insight into how (Fig. 3A-B). We quantify the ground truth proximity of each point on the trajectory to the limit sets using the error coordinates $e_{\rho}$ and $e_{\delta}$, which are the Euclidean distances from the point to the surfaces $\{z:\rho_{1},\rho_{2}=1\}$ and $\{z:y=0\}$, respectively. Geometrically, $e_{\rho}$ and $e_{\delta}$ are the Euclidean distances of the point to the attracting sets $\mathcal{C}$ and $\mathcal{P}$ in Table 1, respectively. When $\gamma_{\rho}$ is large relative to $\gamma_{\delta}$ and the initial state lies off of each subsystem’s attracting set (quantified by the error coordinates, Fig. 3B), the rate at which the distance estimator converges initially outpaces the slower subsystem, reflecting the faster transient. After the fast transient has decayed (indicated temporally by the gray line $\mathcal{C}$ in Fig. 3B and spatially by the gray ’cylindrical’ point cloud in Fig. 3A), the behavior of the distance estimator rapidly switches, and converges at the rate of the slower system. Comparison of the results for systems with and without a limit cycle indicates an effect on the behavior of the distance estimator, with relatively higher values of the projection residual persisting when $\gamma_{\theta}=1$ and $\gamma_{\delta}>\gamma_{\rho}$. This is not unexpected - stabilizing the limit cycle effectively introduces a third time scale set by $|\gamma_{\theta}|$ but does motivate further investigation of VERT for systems with more than two distinct time scales.
@@
Detection of discontinuous or jump phenomena is necessary for inference of systems with non-smooth dynamics, such as legged locomotion or manipulation. The global dynamics of the A-vdP system are characterized by an attracting fixed point, Hamiltonian periodic orbits, or a limit cycle depending on the choice of damping parameters $\mu$. We hypothesize that when $\mu$ is large and the global limit cycle exhibits relaxation oscillations, the ILG nature of VERT will afford sensitivity to not only the globally attracting limit cycle, but also to the local attracting dynamics on the limit cycle.
To test this hypothesis and explore the efficacy of VERT accross dynamical regimes of the A-vdP system, we use VERT to predict attracting sets from simulated data for four values of the bifurcation parameter $\mu$ (Fig. 4). We compare representative predictions from VERT to the ground truth trajectories from the A-vdP system (Fig. 4B). The position coordinate of the anchoring subsystem, $x_{3}$, goes to zero as the trajectory approaches the van der Pol subsystem plane. The acceleration of the van der Pol subsystem, $\dot{x}_{2}$, only approaches zero asymptotically when the attracting set is a globally attracting fixed point ($\mu=-1$). Jump behaviors on the limit cycle are accompanied by a sudden change in the sign of $\dot{x}_{2}$.
@@
VERT Reveals Control Modules in Running Data
@@
Figure 5: Illustrative VERT predictions of running control modules are supported by comparison to empirically validated template models. A. The postural state data input to VERT are presented as trajectories through the space of spatial pairwise distances (and their velocities) between body markers, $z\in\mathbb{R}^{182}$. B. Visualization of the stance phase postural trajectories colored by value of the distance estimator, projected onto the first 3 global principal components (GPCs) of the postural state space. Arrow indicates direction of flow, from impact to lift-off. C. Stance phase averaged time series (with shadowed data traces) of the distance estimator $f(z)$, aligned with vertical ground reaction force (GRF) $F_{z}$. D-E. We use VERT to search for control modules in the Arm-Trunk (D) and Leg (E) postural state projections. We align and compare the averaged distance estimator $f(z)$ to the average vertical GRF time series ($F_{z}$) and kinematics informed by template hypotheses - the trunk and arm pitch angles ($\psi_{t}$ and $\psi_{a}$, respectively, in row D) and the virtual leg length from the COM to heel ($l_{v}$ in row D). The template phase of stance predicted by VERT aligns with the canonical kinematic behavior of the SLIP model [1, 42].
Predictions from the VERT framework benefit from trajectory datasets containing large variation in initial states and rich time-dependent disturbances, as samples on the fiber of the attracting set provide valuable data for estimating the vertical component of the linearized vector field. However, even relatively steady state legged locomotion behaviors contain considerable variability due in part to the presence of sensorimotor noise and latency. Postural kinematic variability is often excluded from behavior analysis during pre-processing, e.g., by projecting trajectories onto reduced latent coordinates such as global principal components to highlight stereotyped features that are useful for downstream tasks such as segmentation and labeling [43]. However, recent experimental and simulation studies on human running show that this intrinsic variability contains critical information for explaining task-level stability mechanisms [44]. Even in less energetic behaviors such as manual grasping, trajectory components excluded by both linear and nonlinear dimensionality reduction contain task-specific structure that could be valuable for understanding sensorimotor control [8].
We test the efficacy of VERT for discovering attracting sets in real postural dynamics data using an open source dataset of whole-body human treadmill kinematics [45] (published in [46]). Bipedal running is an ideal setting for validating the performance of VERT, as interpretable and empirically validated template-anchor models such as the spring loaded inverted pendulum (SLIP) [42] and their characteristic kinematics can be readily extracted from the trajectories for comparison with the predictions from VERT. Because the SLIP template model represents a single virtual leg restricted to the sagittal plane, we focus our analysis on sensors from the right side of the body during the stance phase to facilitate comparison (Fig. 5A).
@@
The VERT predictions for the whole body template are shown in Fig. 5B-C. Visualizing the distance estimator over the spatiotemporal projection of the trajectories onto their leading GPCs (Fig. 5B) reveals a phase of transient postural dynamics immediately following impact, with closest proximity to the template posture occurring during the latter half of stance. Temporal alignment of the distance estimator with the vertical ground reaction force ($F_{z}$) shows that the initial post-impact transient predicted by VERT aligns with the impact transient phase of $F_{z}$ (Fig. 5C, red shaded region), and decays near the transition to the post-transient (SLIP-like) GRF profile. Interestingly, VERT predicts the persistence of a weaker transient following the transition to SLIP-like GRF behavior. This transient decays near the functional transition from energy absorption to propulsion, where the slope of $F_{z}$ changes sign.
@@
These predictions support the longstanding practice of modeling running dynamics as a hybrid limit cycle system. The results of our experiments in the preceding sections suggest that we can take a step further towards identifying the responses of specific appendages and body segments (i.e., control modules) to impact by first projecting the trajectories to the corresponding subspace of the full observation space, and then applying VERT to infer convergence of the projected dynamics to the global attracting set. We explore two control module hypotheses informed by existing template-anchor models of running dynamics: a control module for the leg posture (Fig. 5D), and a control module for the arm and torso posture (Fig. 5E).
@@
The SLIP template and its anchoring in more detailed leg morphologies [47] suggest that the stance leg targets the behavior of a virtual spring. However, empirical measurements of ground reaction forces and COM kinematics show SLIP-like behavior only after an initial transient phase following impact. The distance estimator for the leg projection converges after the $F_{z}$ impact transient (Fig. 5E), with the second transient apparent in the full body projection absent, supporting the hypothesis that the initial postural transient results from stabilization of the SLIP posture. We estimate the length of the ’virtual leg’, $l_{v}$, as the distance from the COM to the right medial ankle marker. Note that only postural variables were used by VERT, so information about the COM position and $l_{v}$ was not directly available to VERT when computing the distance estimator. Convergence of the distance estimator to its sublevel set coincides with the transition to SLIP-like kinematics for $l_{v}$, further supporting that VERT has accurately predicted the transition to the template posture.
@@
Whole-body template models that include arm and trunk dynamics are relatively less studied, in part due to the difficulty of inferring the control target of the more complex multi-body mechanical system. The sublevel set, determined by inspection (Fig. 5D), coincides with the propulsive phase of stance (the phase following the peak of the vertical GRF). Interestingly, the template phase predicted by VERT coincides with the transition to nominally steady state kinematics of the arm posture, quantified by the spatial angle between the proximal segment of the stance arm and the gravity vector ($\psi_{a}$). The trunk-gravity pitch angle ($\psi_{t}$) also exhibits steady state behavior during the predicted template phase, although the onset occurs prior to the convergence of the distance estimator to the sublevel set, suggesting sensitivity primarily to the arm swing dynamics for the arm-trunk projection. Note that both $\psi_{a}$ and $\psi_{t}$ contain privileged information (orientation with respect to gravity) not used by VERT to compute the distance estimator. These results suggest that arm swing dynamics may have an important role for task-level control during the propulsive phase of stance, motivating further study.

### v2

@@ I Introduction
We provide a proof of Theorem 1 in Section 3 of the supplement. The key implication of this result is that for sufficiently small $C_{G}$, the estimator $f$ will be away from 0 together with the distance $|y_{*}|$ to the attractor, in a quantifiable manner, allowing one to ascertain the proximity to it. We remark that the definition, and therefore the estimation and computation of $f$, requires exclusively linear operations on local data in the neighborhood of a point, and does not depend on the split of the coordinate system into $x$ and $y$, despite the fact that the formulation of our estimate does. By computing $f$ for all (or an appropriate subsample) of the trajectory data, and then filtering the $f$-sublevel sets of the full trajectory dataset (geometrically, a point cloud in $\mathbb{R}^{D}$), we can isolate the set of trajectory segments that lie on or quantifiably near any physically relevant attractor $\mathcal{A}$ in the global embedding (observation) space coordinates $\mathbb{R}^{D}$.
@@
We detail algorithms for computing $f(z)$ on measurement data in Section 4 of the supplement. The computation is, at a high level, performed as follows. First we estimate a local Cartesian basis for $T_{z_{*}}\mathcal{M}\simeq T_{x_{*}}\mathcal{A}\oplus\mathbb{R}^{n-a}$ and compute the restriction of the local vector field to this subspace $v_{z_{*}}$. Local PCA, performed on the subsample of points lying within an open Euclidean ball of radius $r$ centered at $z_{*}$, provides a simple approach for computing a local basis of $T_{z_{*}}\mathcal{M}$ such that estimators for the tangent space and intrinsic dimension enjoy strong formal guarantees [35]. The local sample of the vector field $V_{z_{*}}:\mathbb{R}^{D}\rightarrow T\mathbb{R}^{D}$ can be approximated from finite differences. The local representation of the ($n$-dimensional) $v_{z_{*}}$ can then be found by orthogonal projection of the ($D$-dimensional) $V_{z_{*}}$ onto the vielbein basis. The Jacobian $H(z_{*})$ can then be estimated by centering the local vector field data and using linear estimation techniques; for experiments in the following sections, we use least squares regression with the standard MSE residual.
Figure 2: Application of VERT to the A-HM model illustrates the mechanisms laid out in Fig. 3 that underlie the hypotheses in Table 1: the distance estimator is sensitive to the locally dominant dynamics at the sampled point. A. Hierarchical composition of the A-HM (Eqns. 2 and 3). The cross product system of two limit cycles produces a limit torus, whose product with a higher dimensional stable spiral generates qualitatively distinct mechanisms for anchoring the global attractor. B. Visualization of VERT estimator ($f$) in the error coordinate plane. Observe the initial alignment of the $f$-level sets with the error coordinate of the stronger subsystem. Upon reaching an intermediate attractor (at vertical and horizontal axes), $f$ continues to decay more slowly to its global minimum. C. Sensitivity plots of $f$ for different relative subsystem gains. When $\gamma_{\rho}=\gamma_{\delta}$, $f$ varies linearly with both error coordinates. When one gain is larger, $f$ shows heightened sensitivity to the stronger system.
Figure 3: Parameter sweep experiments illustrate sensitivity of the VERT estimator to attractor hierarchies. See Table 1 for hypothesis definitions A. When $\gamma_{\theta}=0$, the limit cycle $\mathcal{S}$ does not exist, and the outputs of VERT accordingly support rejection of $H_{0}$. $H_{1-3}$ are supported, as the residual decreases with both $(\gamma_{\rho},\gamma_{\delta})$ for $\mathcal{T}$, $\gamma_{\delta}$ for $\mathcal{P}$, and $\gamma_{\rho}$ for $\mathcal{C}$. B. When $\gamma_{\theta}=1$, $\mathcal{S}$ exists and is detected. $H_{0-3}$ are supported, with similar behavior to the $\gamma_{\theta}$ for the intermediate submanifolds.
@@ Results
Detecting the presence of subprocesses evolving on distinct time scales is critical for model inference and interpretation of systems constructed by parallel composition of stable subsystems. We hypothesize that the ILG nature of the VERT pipeline affords sensitivity to both global and intermediate attracting sets through a detectable change in the behavior of the distance estimator along a trajectory as it transitions between them. We construct a model system inspired by dynamic movement primitives, the anchored Hopf model (A-HM). The A-HM embeds a pair of coupled Hopf-type limit cycles in a higher dimensional Euclidean space via a hyperbolic linear flow. The A-HM template-anchor system is easiest to render in polar coordinates. The coupled limit cycle subsystem is:
$split&\dot{\rho}_{i}=\gamma_{\rho}(1-\rho_{i})\rho_{i}\\ &\dot{\theta}_{i}=\omega+\gamma_{\theta}\sin(\theta_{i}-\theta_{j})$ (2)
where $i,j\in\{1,2\}$. We use the variable $x\in\mathbb{R}^{D}$ to refer to the corresponding Euclidean coordinate. Observe that the dynamics of the phase variables decouples from the amplitudes $\rho$, and has the structure of a Kuramoto oscillator system. The system entrains to a global limit cycle so long as $\gamma_{\theta}>0$. The auxiliary embedding subsystem has the form:
$\dot{y}=Ay$ (3)
with a Hurwitz $A$, and $y=(y_{5},\ldots,y_{D})$ is the vector of the auxiliary variables. The state vector for the full system composed of Eqs. 2 & 3 is thus $(z_{1},...,z_{D})=(x_{1},...,x_{4},y_{5},...,y_{D})\in\mathbb{R}^{D}$. The rate of convergence to the anchored plane at $y=0$ is set by the real part of the eigenvalues $\sigma$ of $A$, $\Re\sigma=-\gamma_{\delta}$.
We test the sensitivity hypotheses elaborated in Table 1 using simulated trajectories from the A-HM, while systematically varying the control parameters $\gamma_{\rho}$ and $\gamma_{\delta}$, which determine the rate of convergence for the two subsystems. We define the error coordinates $e_{\rho}$ and $e_{\delta}$ as the normalized Euclidean distance from a point on the trajectory to the limit torus of Eq. 2 and limit point of Eq. 3, respectively. These error coordinates provide ground truths for the global fiberwise distances that VERT estimates. The qualitative behavior of $f$ across the $(\gamma_{\rho},\gamma_{\delta}$) plane is illustrated in Fig. 2. In these experiments we set $\gamma_{\theta}=0$ without loss of generality, so that the limit torus is the ultimate global attracting submanifold. Trajectories can either converge directly to the limit torus $\mathbb{T}^{2}$, or first approach either the plane of the limit cycle subsystem ($\mathcal{P}\cong\mathbb{R}^{4}$) when $\gamma_{\delta}>>\gamma_{\rho}$ or the space $\mathcal{C}\cong\mathbb{T}^{2}\times\mathbb{R}^{16}$ when $\gamma_{\rho}>>\gamma_{\delta}$, before converging to the limit torus. We quantify proximity with the normalized error coordinate $e_{\star}$, and observe empirically that when $\gamma_{\delta}>\gamma_{\rho}$, $\frac{\partial f}{\partial e_{\delta}}>\frac{\partial f}{\partial e_{\rho}}$ in the near-field of the attractor and vice-versa, indicating the switch in sensitivity of the estimator when the relative magnitudes of the gains are permuted. A table summarizing these detection hypotheses and a visual rendering of the relationships between trajectories and their intermediate attracting submanifolds is included in Section 5B of the Supplement.
Subsystem Gain Anchoring Hierarchy VERT Detection Hypothesis
$(\mathbf{H_{0}})$ $\gamma_{\theta}\gg 0$ $z_{i}\in\mathbb{R}^{d}\rightarrow\cdot\cdot\cdot\rightarrow\mathcal{S}\ni z_{f}$ Detect $\mathcal{S}$
Table 1: Table of hypothesized anchoring mechanisms ($\mathbf{H_{0-3}}$) for the A-HM system (Eq. 2, 3). The choices of subsystems gains modulate which intermediate attracting submanifolds a trajectory visits while approaching the global invariant submanifold. We study the sensitivity of VERT to such hierarchical systems by testing the detection hypotheses in the right column.
To develop a systematic test of the detection hypotheses, we use the continuous output $f$ to construct a binary classifier that predicts the ”template” set: the trajectory segments that lie sufficiently near the globally invariant attracting set. In the context of our theory, the template set can be viewed as a finite approximation to the zero set $f^{-1}(0)$. We then quantify the quality of the template set estimate as the MSE (mean squared error) projection residual, the mean over orthogonal Euclidean projection distances from each point in the template set to the true submanifold. Lower values of the residual indicate that the predicted template set more closely conforms to the hypothesized submanifold, while relatively higher values indicate that the identified template set contains a larger proportion of samples distant from the target attracting submanifold. The results are summarized in Fig. 3 for experiments where the lowest level attractor is either the limit torus ($\gamma_{\theta}=0$) or the limit cycle ($\gamma_{\theta}=1$). When one of the anchoring gains $(\gamma_{\rho},\gamma_{\delta})$ is larger than the other, relatively lower magnitudes of the residual are achieved for the corresponding submanifold, with relatively higher residual for the submanifold corresponding to the weaker gain. We note that there are qualitative differences in the results over the $(\gamma_{\delta},\gamma_{\rho})$ parameter space between the experiments with and without a global limit cycle. We note that varying $\gamma_{\theta}$ effectively introduces a third time scale, which in turn can influence the behavior of $f$. However, the behavior of the VERT predictions over the $(\gamma_{\rho},\gamma_{\delta})$ parameter space is empirically consistent for both settings of $\gamma_{\theta}$.
@@
Detection of discontinuous or jump phenomena is necessary for inference of systems with non-smooth dynamics, such as legged locomotion or manipulation. We analyze the behavior of VERT on simulated data from a continuous system with an emergent jump behavior, termed the anchored van der Pol system (A-vdP). A-vdP is a van der Pol type oscillator anchored in $\mathbb{R}^{4}$ by an overdamped harmonic oscillator. In Cartesian coordinates:
$split&\dot{x}_{1}=x_{2}\\ &\dot{x}_{2}=-\sin(x_{1})+\mu(\cos^{2}(x_{1})-0.5)x_{2}\\ &\dot{x}_{3}=x_{4}\\ &\dot{x}_{4}=-x_{3}-bx_{4}$ (4)
Where $b$ is the damping coefficient for the harmonic oscillator. Depending on the choice of the continuous parameter $\mu$, the global dynamics of the A-vdP system are characterized by an attracting fixed point, Hamiltonian periodic orbits, or a limit cycle with relaxation oscillations. When $\mu$ is sufficiently large, the restriction dynamics to the limit cycle are interpreted as an effectively hybrid dynamical system with fast jumps between relatively slow hyperbolic segments [55]. In principle, the global limit cycle could be discovered from data as the locus of fixed points of the Poincare section, but this analysis would not provide insight into the jump behavior or locally hyperbolic nature of the flow on the limit cycle itself. We hypothesize that when $\mu$ is large enough that the limit cycle exhibits jump phenomena, the ILG nature of VERT will allow identification of the locally hyperbolic subdomains.
To test this hypothesis and explore the efficacy of VERT across dynamical regimes of the A-vdP system, we use VERT to predict attracting sets from simulated data for four values of the bifurcation parameter $\mu$ (Fig. 4). We compare representative predictions from VERT to the ground truth trajectories from the A-vdP system (Fig. 4B). The position coordinate of the anchoring subsystem, $x_{3}$, goes to zero as the trajectory approaches the van der Pol subsystem plane. The acceleration of the van der Pol subsystem, $\dot{x}_{2}$, only approaches zero asymptotically when the attracting set is a globally attracting fixed point ($\mu=-1$). Jump behaviors on the limit cycle are accompanied by a sudden change in the sign of $\dot{x}_{2}$.
@@
VERT Reveals Candidate Control Modules in Human Running Data
@@
Figure 5: Illustrative VERT predictions of running control modules are supported by comparison to empirically validated template models based on data withheld from the VERT pipeline. Blue traces show the VERT distance estimate. A. The postural state data input to VERT are presented as trajectories through the space of spatial pairwise distances (and their velocities) between body markers, $z\in\mathbb{R}^{182}$. B. Visualization of the stance phase postural trajectories colored by value of the distance estimator, projected onto the first 3 global principal components (GPCs) of the postural state space. Arrow indicates direction of flow, from impact to lift-off. C. Stance phase averaged time series (with shadowed data traces) of the distance estimator $f(z)$ (blue trace), aligned with vertical ground reaction force (GRF), $F_{z}$ (red trace, withheld from VERT). D-E. We use VERT to search for control modules in the Arm-Trunk (D) and Leg (E) postural state based only upon projections of the purely kinematic data (visualized in the lower plot of A). As a validation experiment probing VERT’s ability to find postures carrying previously identified template dynamics (E), and to suggest novel postures that might carry new templates (D), we align and compare the averaged distance estimator $f(z)$ to the average vertical GRF time series ($F_{z}$) and kinematics informed by the previously known template hypotheses - the trunk and arm pitch angles ($\psi_{t}$ and $\psi_{a}$, respectively, in row D) and the virtual leg length from the COM to heel ($l_{v}$ in row D). The template phase of stance predicted by VERT’s proposal of a stabilized posture submanifold (the blue shaded interval that marks the kinematic data at which $f(z)$ vanishes) aligns with the canonical dynamic behavior of the SLIP model [21, 29].
The presence of low dimensional spring loaded inverted pendulum (SLIP) dynamics embedded in the high dimensional postures of running, trotting, and hopping legged animals spanning mass and length scales is a foundational principle of legged locomotion [4]. Observation and identification of SLIP dynamics in data has required model-based analysis of dynamics variables including ground reaction force (GRF) and global center of mass position. The ability to detect the presence of such template dynamics from kinematics data alone would enable the discovery of new classes of dynamical locomotion templates without imposing the biases inherent in fitting a pre-determined model. However, such a procedure has remained elusive. Dynamical systems theory suggests that stability of the low dimensional dynamics requires stability of a characteristic low dimensional posture [29].
We use publicly available motion capture data of human treadmill running [39] (published in [40]) as a validation experiment to test VERT’s capacity to detect template dynamics from postural kinematics measurements alone. The canonical latent variables of the SLIP embedding are the state variables of an extensible ”virtual leg” extending from the center of mass to the contact foot [29, 43], and the characteristic stance dynamics of an inverted spring mass system are observable from privileged measurements - GRF and COM position - available in the dataset, but withheld from VERT.
Predictions from the VERT framework benefit from trajectory datasets containing large variation in initial states and rich time-dependent disturbances, as the primary object of estimation is the pointwise restriction of the flow to the fibers of the invariant set, rather than the flow on the invariant set itself. Treadmill running is a relatively steady state behavior, with perturbing forces arising primarily during the impulsive dynamics of foot-ground contact. However, even relatively steady state legged locomotion behaviors contain considerable variability due in part to the presence of sensorimotor noise and latency. In quantitative analyses of behavior, it is standard to reduce postural kinematic variability by projection to latent subspaces (e.g., global principle components) to isolate essential stereotyped features that are useful for downstream tasks such as segmentation and labeling [6]. Our analysis with VERT (Fig. 5) demonstrates a complementary perspective: the transient ”off-manifold” dynamics of the trajectory ensemble contain valuable information about the mechanisms that stabilize the observed pattern of behavior.
@@
Because the SLIP template model represents a single virtual leg restricted to the sagittal plane, we focus our analysis on sensors from the right side of the body during the stance phase to facilitate comparison (Fig. 5A). The VERT predictions for the whole body template are shown in Fig. 5B-C. Visualizing the distance estimator over the spatiotemporal projection of the trajectories onto their leading GPCs (Fig. 5B) reveals a phase of transient postural dynamics immediately following impact, with closest proximity to the template posture occurring during the latter half of stance. Temporal alignment of the distance estimator with the vertical ground reaction force ($F_{z}$), withheld from VERT, shows that the initial post-impact transient predicted by VERT aligns with the impact transient phase of $F_{z}$ (Fig. 5C, red shaded region), and decays near the transition to the post-transient (SLIP-like) GRF profile. Interestingly, VERT predicts the persistence of a weaker transient following the transition to SLIP-like GRF behavior. This transient decays near the functional transition from energy absorption to propulsion, where the slope of $F_{z}$ changes sign. Low values of the distance estimator thus align with SLIP template behavior inferred from the privileged observables (vertical GRF and virtual leg length).
@@
These predictions support the longstanding practice of modeling running dynamics as a hybrid limit cycle system. The results of our experiments in the preceding sections suggest that we can take a step further towards identifying the responses of specific appendages and body segments (i.e., control modules) to impact by first projecting the trajectories to the corresponding subspace of the full observation space, and then applying VERT to infer convergence of the projected dynamics to the attracting posture submanifold presumed tocarry some template dynamics. We explore two control module hypotheses. The first is informed by existing template-anchor models of running dynamics: a control module for the leg posture expected to carry the SLIP template (Fig. 5E). The second posits a control module for the arm and torso posture that the template/anchor framework would presum to carry some template which is, nevertheless, presently unknown (Fig. 5D).
@@
Empirical measurements of ground reaction forces and COM kinematics withheld from VERT show SLIP-like behavior only after an initial transient phase following impact. VERT’s distance estimator for the leg projection converges after the $F_{z}$ impact transient (Fig. 5E), with the second transient apparent in the full body projection absent, supporting the hypothesis that the initial postural transient results from stabilization of the SLIP posture. We estimate the length of the ’virtual leg’, $l_{v}$, as the distance from the COM to the right medial ankle marker. We again emphasize that only postural variables were used by VERT, so information about the COM position and $l_{v}$ was not directly available to VERT when computing the distance estimator. Convergence of the distance estimator to its near-zero values444ff takes its values in the co-domain of real numbers, whereas its level sets are comprised of points in its domain of data samples. identifies a sublevel set of data samples that coincides precisely with the transition to SLIP-like kinematics for $l_{v}$, supporting the validation hypothesis that VERT has accurately predicted the transition to the posture carrying the template.
@@
Whole-body bipedal template running models that include arm and trunk dynamics are relatively less studied, in part due to the difficulty of inferring the control target of the more complex multi-body mechanical system. However, we can observe, again by comparison with the additional GRF data withheld from VERT, that its distance estimator’s near-zero sublevel set of arm and trunk projected data samples determined by inspection (Fig. 5D), coincides with the propulsive phase of stance (the phase following the peak of the vertical GRF). Interestingly, the template phase predicted by VERT coincides with the transition to nominally steady state kinematics of the arm posture, quantified by the spatial angle between the proximal segment of the stance arm and the gravity vector ($\psi_{a}$). The trunk-gravity pitch angle ($\psi_{t}$) also exhibits steady state behavior during the predicted template phase, although the onset occurs prior to the convergence of the distance estimator to the sublevel set, suggesting sensitivity primarily to the arm swing dynamics for the arm-trunk projection. Note that both $\psi_{a}$ and $\psi_{t}$ contain privileged information (orientation with respect to gravity) not used by VERT to compute the distance estimator. These results suggest that arm swing dynamics may have an important role for task-level control during the propulsive phase of stance, highlighting a promising direction for future study.
Finally, we note that these results complement recent experimental and simulation studies on human running showing that intrinsic variability contains critical information for explaining task-level stability mechanisms [49]. Even in less energetic behaviors such as manual grasping, trajectory components excluded by both linear and nonlinear dimensionality reduction contain task-specific structure that could be valuable for understanding sensorimotor control [60].